# Your Paper Title Here

### v1

@@ Abstract
We here investigate whether socio-economic indicators, such as household wealth, leave recoverable informational imprints in both satellite imagery (capturing physical features like buildings and roads) and Internet-sourced text (reflecting historical, cultural, and economic narratives of neighborhoods). Using Demographic and Health Survey (DHS) data from African neighborhoods (clusters), we pair high-resolution Landsat images with textual descriptions generated by large language models (LLMs) conditioned on location and year, as well as text retrieved by an LLM-driven AI Search Agent from web sources. We develop a multimodal framework that predicts household wealth (measured by the International Wealth Index (IWI)) through five pipelines: (i) a vision model on satellite images, (ii) an LLM using only location and year, (iii) an AI agent that searches and synthesizes web text, (iv) a joint image-text encoder, and (v) an ensemble of all signals. Our framework yields three contributions. First, evaluations show that fusing vision and agent/LLM-generated text outperforms vision-only baselines in wealth prediction (e.g., $R^{2}=0.77$ vs. 0.630.630.63 on out-of-sample splits), with LLM-internal knowledge (artificial neural memory) proving surprisingly more effective than agent-retrieved text, improving robustness to out-of-country and out-of-time generalization. Second, we find partial representational convergence: fused embeddings from vision and language modalities correlate moderately (median cosine similarity across modalities of about 0.60 after alignment), suggesting a shared latent code of material well-being while retaining complementary details, broadly consistent with the Platonic Representation Hypothesis. Although the superior performance of LLM-only text over agent-retrieved data challenges our Agent-Induced Novelty Hypothesis, modest gains from combining agent data in some splits offer weak support for the idea that agent-gathered information—emerging from dynamic interaction with the Internet—introduces a degree of unique representational structures not fully captured by static LLM knowledge. Third, we release a large-scale multimodal dataset comprising approximately 60,000 DHS clusters, each linked to satellite images, LLM-generated descriptions, and associated texts retrieved by AI agents.
@@ 2 Problem Setup & Related Work
Our work builds on the framework outlined in the Introduction (Figure 1), where the latent poverty/wealth index $Y_{i}$ causally influences both satellite imagery $\mathbf{M}_{i}$ and textual data $\mathbf{W}_{i}$, assuming conditional independence between modalities given $Y_{i}$. This structure motivates our investigation into whether vision and language encoders converge toward a shared “Platonic” representation of material well-being, as inspired by the Platonic Representation Hypothesis as proposed by Huh et al. (2024). While this hypothesis has been examined in general vision-language models (Radford et al., 2021), its application to specific scientific domains, such as economic development and poverty mapping—particularly in contexts where data modalities are gathered endogenously by AI agents—is novel.
@@ 3 Data & Methods
The resulting combined corpus—images, LLM texts, agent traces, with IWI labels, dubbed IWI-Africa-Multimodal—will be released on Hugging Face. [^1] To link with the full DHS data, users must register with the DHS Program and agree to its privacy and data use policies, as outlined at https://dhsprogram.com. See Table 1 for two illustrative AI Search Agent traces.
@@
Here, embeddings refer to dense vector representations of input data (e.g., images or text) learned by neural networks, which capture semantic or visual features in a compact, latent space suitable for downstream tasks like regression. In our pipelines, we distinguish between (1) frozen embeddings, which are pre-computed from pre-trained models—e.g., OpenAI’s text-embedding-3-small [abbreviated as OAI] or sentence-transformers/all-mpnet-base-v2, abbreviated as [MPNet]—and not updated during training to preserve general knowledge, and (2) fine-tuned embeddings, which are adjusted on our dataset to adapt to poverty-specific patterns. This distinction allows us to evaluate the trade-offs between leveraging off-the-shelf representations and task-specific optimization.
@@
  • Out-of-Country (OOC): The model is trained on clusters from a subset of countries (a random 80% subset of the data) and evaluated on held-out countries not present in the training set (the remaining 20%). This test evaluates cross-border generalization and the model’s ability to transfer knowledge beyond national boundaries.
@@ 4 Results
Figure 4: Spatial distribution of the difference in prediction residuals between CV-only and NMR+CV. Blue indicates regions where the baseline CV model outperforms NMR+CV. Red indicates regions where NMR+CV outperforms CV. Hexagons summarize mean residual differences across locations.
@@
Our evaluation reveals significant performance gains when combining multi-modal signals for poverty prediction, as shown in Figure 3 (full results in Table A.I.1). The NMR+CV approach using Llama-4-Maverick with OpenAI embeddings achieves the highest performance across all evaluation strategies, with an $R^{2}$ of 0.765 on the random split. This represents a substantial improvement over the best single-modality approach (NMR alone at $R^{2}$ = 0.668) and a significant advancement over the CV-only baseline ($R^{2}$ = 0.634). Our CV-only baseline outperforms the shallow baseline method by Pettersson et al. ($R^{2}$ = 0.60 on OOC split), despite using only daytime satellite imagery (instead of both daytime and nighttime images, which is known to boost performance). This demonstrates the effectiveness of our pipeline and suggests that our multimodal approach can achieve comparable or better results with fewer data sources. The performance gains are consistent across evaluation strategies, with the combined model maintaining strong performance even under OOC and out-of-time (OOT) splits. It is notable that the ASA searchers alone perform below the CV-only benchmark, possibly indicating that finding relevant online information is a noisy process.
@@
To further contextualize performance, we conducted an extensive analysis of different model architectures and data sources, building on the textual representations defined earlier: $\mathbf{W}_{i}^{\text{Trace}}$ (raw agent search traces), $\mathbf{W}_{i}^{\text{Summary}}$ (agent-synthesized summaries), and $\mathbf{W}_{i}^{\text{Desc}}$ (LLM-generated descriptions). Here, CleanedTraces refers to all agent-crawled text data ($\mathbf{W}_{i}^{\text{Trace}}$) without filtering; Wikipedia is the subset from Wikipedia sources; JustificationOnly includes only the agent’s justification of IWI prediction text; and JustificationPrediction combines justification with the agent’s scalar prediction. The results reveal that the CleanedTraces approach consistently outperforms other subset text sources, including Wikipedia, JustificationOnly, and JustificationPrediction. This finding suggests that the full breadth of agent-generated context provides richer signals for poverty prediction.
@@
Our analysis of different embedding models reveals that text-embedding-3-small (OAI) consistently outperforms the fine-tuned MPNet model (sentence-transformers/all-mpnet-base-v2) across all evaluation strategies (see Table A.I.1 and Figure 3 for detailed results). For instance, in the NMR+CV pipeline under random splits, OAI achieves an $R^{2}$ of 0.765 compared to MPNet’s 0.747. MPNet cannot match the performance of the frozen-weights OAI model. This suggests that the pretraining and contextual understanding captured in OAI provide a significant advantage for poverty prediction tasks over a smaller model fine-tuned on our dataset.
@@
To quantify spatial variability in gains from multimodality, Figure 4 plots the following difference in absolute residuals, $(\lvert Y_{i}-\hat{Y}_{i}^{\text{CV}}\rvert-\lvert Y_{i}-\hat{Y_{i}}^{\textrm{Best}}\rvert)$. This value is negative (blue) when the CV baseline outperforms the best multi-modal approach; it is positive (red) otherwise. We benchmark here the best-performing joint model against the computer-vision model, as the CV approach is dominant in the ML literature on poverty prediction.
@@
To complement this spatial analysis, we also analyzed where the temporal improvements are occurring, using the same metric as for the Africa-wide spatial evaluation. Figure 5 shows that the improvements are consistent over time, where the multi-modal model outperforms the CV-only model. Most of the residual improvements are occurring in the 1990s. Interestingly, this is the period when satellite images are most scarce, with an average pixel availability of about three-quarters across the continent, from 1990 to 2000 Jerzak et al. (2023).
Figure 5: Absolute mean residual difference across years for the best system combining NMR+ASA+CV vs CV-only baseline.
@@
One of the most notable findings from our experiments is that the LLM-only NMR approach consistently outperforms the ASA approaches. For instance, NMR with Llama-4-Maverick achieves $R^{2}$ = 0.668 on Random Split, while the best ASA approach (CleanedTraces) only achieves $R^{2}$ = 0.606, indicating that the LLM’s internal knowledge is more predictive than agent-retrieved text. This contradicts our initial expectation that agent-retrieved data would provide more valuable live context and instead aligns with findings in Gema et al. (2025), which show reduced performance in language tasks as context (e.g., via reasoning chains) becomes overly large.
@@
Figure 6: Histogram of cosine similarities between NMR and CV embeddings, after alignment through canonical correlation analysis (first component), for matched DHS clusters.
To statistically validate the observed alignment in Figure 6, we performed a one-sample $t$-test on the cosine similarities between NMR and CV embeddings for matching clusters, testing the null hypothesis that their mean is zero (with empirical $\sigma\approx 0.11$ derived from random pairings). The test produced a $t$-statistic of 312.96 and a $p$-value $<1\times 10^{-10}$, decisively rejecting the null and confirming significant positive similarities. This evidence supports partial representational convergence across modalities, consistent with the Platonic Representation Hypothesis, as vision and text embeddings share meaningful semantic structures beyond random chance while contributing complementary details to poverty prediction.
@@
Figure 7: Latent representation similarity matrix of matching embeddings from NMR and CV, based on cosine similarity after alignment through canonical correlation analysis (first component). Embeddings are sorted by latitude to allow regional interpretability. See Figure A.I.3 for ASA/CV comparison.
@@
To this end, we compute a similarity matrix between NMR-based and CV-based embeddings using cosine similarity after alignment through canonical correlation analysis (Weenink, 2003), with rows/columns sorted by latitude for regional interpretability (Figure 7). The resulting matrix reveals localized clusters of higher similarity, suggesting that embeddings from geographically proximate regions along similar latitudes exhibit stronger alignment. This spatial coherence indicates that geographically proximate areas exhibit systematically stronger cross-modal alignment.
@@ 5 Discussion & Conclusion
Limitations of our study include reliance on DHS clusters, which may introduce sampling biases by excluding remote areas from sampling (Kakooei and Daoud, 2024). The search agent may also retrieve post-treatment information from the web, potentially biasing causal analyses. Additionally, the computational demands of agent-driven pipelines at continental scales limit scalability for global or time-series applications.
Future work could integrate causal inference into multimodal socio-economic AI while addressing text-specific challenges (Grimmer et al., 2022; Daoud et al., 2022; Pieuchon et al., 2024)—such as improved detection and control of post-treatment bias in agent representations—to further test the Platonic Representation and Agent-Induced Novelty Hypotheses. $\otimes$

### v2

# Platonic Representations for Poverty Mapping: Unified Vision–Language Codes or Agent–Induced Novelty?

@@ Abstract
We here investigate whether socio-economic indicators, such as household wealth, leave recoverable informational imprints in both satellite imagery (capturing physical features like buildings and roads) and Internet-sourced text (reflecting historical, cultural, and economic narratives of neighborhoods). Using Demographic and Health Survey (DHS) data from African neighborhoods (clusters), we pair high-resolution Landsat images with textual descriptions generated by large language models (LLMs) conditioned on location and year, as well as text retrieved by an LLM-driven AI Search Agent from web sources. We develop a multimodal framework that predicts household wealth (measured by the International Wealth Index (IWI)) through five pipelines: (i) a vision model on satellite images, (ii) an LLM using only location and year, (iii) an AI agent that searches and synthesizes web text, (iv) a joint image-text encoder, and (v) an ensemble of all signals. Our framework yields three contributions. First, evaluations show that fusing vision and agent/LLM-generated text outperforms vision-only baselines in wealth prediction (e.g., $R^{2}=0.77$ vs. $0.63$ on out-of-sample splits), with LLM-internal knowledge (artificial neural memory) proving surprisingly more effective than agent-retrieved text, improving robustness to out-of-country and out-of-time generalization. Second, we find partial representational convergence: fused embeddings from vision and language modalities correlate moderately (median cosine similarity across modalities of about 0.60 after alignment), suggesting a shared latent code of material well-being while retaining complementary details, broadly consistent with the Platonic Representation Hypothesis. Although the superior performance of LLM-only text over agent-retrieved data challenges our Agent-Induced Novelty Hypothesis, modest gains from combining agent data in some splits offer weak support for the idea that agent-gathered information—emerging from dynamic interaction with the Internet—introduces a degree of unique representational structures not fully captured by static LLM knowledge. Third, we release a large-scale multimodal dataset comprising approximately 60,000 DHS clusters, each linked to satellite images, LLM-generated descriptions, and associated texts retrieved by AI agents.
@@ 2 Problem Setup & Related Work
Our work builds on the framework outlined in the Introduction (Figure 1), where the latent poverty/wealth index $Y_{i}$ causally influences both satellite imagery $\mathbf{M}_{i}$ and textual data $\mathbf{W}_{i}$, assuming conditional independence between modalities given $Y_{i}$. This structure motivates our investigation into whether vision and language encoders converge toward a shared “Platonic” representation of material well-being, as inspired by the Platonic Representation Hypothesis as proposed by Huh et al. (2024). While this hypothesis has been examined in general vision-language models (Radford et al., 2021a), its application to specific scientific domains, such as economic development and poverty mapping—particularly in contexts where data modalities are gathered endogenously by AI agents—is novel.
@@ 3 Data & Methods
The resulting combined corpus—images, LLM texts, agent traces, with IWI labels, dubbed IWI-Africa-Multimodal—will be released on Hugging Face. [^1] To link with the full DHS data, users must register with the DHS Program and agree to its privacy and data use policies, as outlined at https://dhsprogram.com. See Table 1 for two illustrative AI Search Agent traces. Table 2 displays the most commonly found URLs in the AI search.
@@
Table 2: Distribution of primary external text sources retrieved by the AI Search Agent across the full IWI-Africa-Multimodal dataset. Wikipedia dominates the retrieved traces, reflecting its role as the most comprehensive open-knowledge repository for African localities.
Domain Cluster Count
wikipedia.org 66,998
mapcarta.com 28,337
researchgate.net 26,375
citypopulation.de 18,705
academia.edu 16,333
city-facts.com 12,395
opendataforafrica.org 12,062
worldbank.org 10,983
@@
Here, embeddings refer to dense vector representations of input data (e.g., images or text) learned by neural networks, which capture semantic or visual features in a compact, latent space suitable for downstream tasks like regression. In our pipelines, we distinguish between (1) frozen embeddings, which are pre-computed from pre-trained models—e.g., OpenAI’s text-embedding-3-small [abbreviated as OAI] or sentence-transformers/all-mpnet-base-v2, [abbreviated as MPNet]—and not updated during training to preserve general knowledge, and (2) fine-tuned embeddings, which are adjusted on our dataset to adapt to poverty-specific patterns. This distinction allows us to evaluate the trade-offs between leveraging off-the-shelf representations and task-specific optimization.
@@
  • Out-of-Country (OOC): The model is trained on clusters from a subset of countries (a random 80% subset) and evaluated on held-out countries not present in the training set (the remaining 20%). This test evaluates cross-border generalization and the model’s ability to transfer knowledge beyond national boundaries.
@@ 4 Results
Figure 4: Spatial distribution of the difference in prediction residuals between CV-only and NMR+CV. Blue indicates regions where the baseline CV model outperforms NMR+CV. Red indicates regions where NMR+CV outperforms CV. Hexagons summarize values across locations.
@@
Our evaluation reveals significant performance gains when combining multi-modal signals for poverty prediction, as shown in Figure 3 (full results in Table A.I.1, with uncertainty estimates in Table A.I.2). The NMR+CV approach using Llama-4-Maverick with OpenAI embeddings achieves the highest performance across all evaluation strategies, with an $R^{2}$ of 0.765 on the random split. This represents a substantial improvement over the best single-modality approach (NMR alone at $R^{2}$ = 0.668) and a significant advancement over the CV-only baseline ($R^{2}$ = 0.634). Our CV-only baseline outperforms the shallow baseline method by Pettersson et al. ($R^{2}$ = 0.60 on OOC split), despite using only daytime satellite imagery (instead of both daytime and nighttime images, which is known to boost performance). This demonstrates the effectiveness of the approach here and suggests that the multimodal approach can achieve comparable or better results with fewer data sources. The performance gains are consistent across evaluation strategies, with the combined model maintaining strong performance even under OOC and out-of-time (OOT) splits. It is notable that the ASA searchers alone perform below the CV-only benchmark, possibly indicating that finding relevant online information is a noisy process.
@@
To further contextualize performance, we conducted an extensive analysis of different model architectures and data sources, building on the textual representations defined earlier: $\mathbf{W}_{i}^{\text{Trace}}$ (raw agent search traces), $\mathbf{W}_{i}^{\text{Summary}}$ (agent-synthesized summaries), and $\mathbf{W}_{i}^{\text{Desc}}$ (LLM-generated descriptions). Here, CleanedTraces refers to all agent-crawled text data ($\mathbf{W}_{i}^{\text{Trace}}$) without filtering; Wikipedia is the subset from Wikipedia sources; JustificationOnly includes only the agent’s justification of IWI prediction text; and JustificationPrediction combines justification with the agent’s scalar prediction. The results reveal that the CleanedTraces approach consistently outperforms all other subset text sources. This suggests that the full breadth of agent-generated context provides richer signals for poverty prediction.
@@
Our analysis of different embedding models reveals that text-embedding-3-small (OAI) consistently outperforms the fine-tuned MPNet model (all-mpnet-base-v2) across all evaluation strategies (see Table A.I.1 and Figure 3 for detailed results). For instance, in the NMR+CV pipeline under random splits, OAI achieves an $R^{2}$ of 0.765 compared to MPNet’s 0.747. MPNet cannot match the performance of the frozen-weights OAI model. This suggests that the pretraining and contextual understanding captured in OAI provide a significant advantage for poverty prediction tasks over a smaller model fine-tuned on our dataset.
@@
[…]
@@
To complement this spatial analysis, we also analyzed where the temporal improvements are occurring, using the same metric as for the Africa-wide spatial evaluation. Figure 5 shows that the improvements are consistent over time, where the multi-modal model outperforms the CV-only model. Most residual improvements occurred in the 1990s. Interestingly, this is the period when satellite images are most scarce, with an average pixel availability of about three-quarters across the continent, from 1990 to 2000 Jerzak et al. (2023).
Figure 5: Absolute mean residual difference across years for the best system: NMR+ASA+CV vs. CV-only baseline.
@@
One of the most notable findings from our experiments is that the LLM-only NMR approach consistently outperforms the ASA approaches. For example, NMR with Llama-4-Maverick achieves up to $R^{2}$ = 0.668, while the best ASA approach (CleanedTraces) only achieves $R^{2}$ = 0.606, indicating that the internal knowledge of LLM is more predictive than agent-retrieved text. This contradicts our initial expectation that agent-retrieved data would provide more valuable live context and instead aligns with findings in Gema et al. (2025), which show reduced performance in language tasks as context (e.g., via reasoning chains) becomes overly large.
@@
Figure 6: Histogram of cosine similarities between NMR and CV embeddings, after alignment through canonical correlation (first component), for matched DHS clusters.
To statistically validate the observed alignment in Figure 6, we performed a one-sample $t$-test on the cosine similarities between NMR and CV embeddings for matching clusters, testing the null hypothesis that their mean is zero (with empirical $\sigma\approx 0.11$ derived from random pairings). The test produced a $t$-statistic of 312.96 and a $p$-value $1\mathrm{e}{-10}$, confirming significant positive similarities. This evidence supports partial representational convergence across modalities, consistent with the Platonic Representation Hypothesis, as vision and text embeddings share meaningful semantic structures beyond random chance while contributing complementary details to poverty prediction.
Our median cosine similarity of  0.60 between Earth Observation (EO) and LLM embeddings reflects moderate convergence in multimodal representations, falling within the typical range of 0.15–0.80 reported in (satellite) imagery and vision-language alignment literature (Dhakal et al., 2024; Radford et al., 2021b), again supporting Platonic Representation Hypothesis over strong Novelty.
@@
Figure 7: Latent representation similarity matrix of matching embeddings from NMR and CV, based on cosine similarity after alignment through canonical correlation. Embeddings are sorted by latitude to allow regional interpretability. See Figure A.I.3 for ASA/CV comparison.
@@
To this end, we compute a similarity matrix between NMR- and CV-based embeddings using cosine similarity after alignment through canonical correlation analysis (CCA) (Weenink, 2003), with rows/columns sorted by latitude for regional interpretability (Figure 7). The resulting matrix reveals localized clusters of higher similarity, suggesting that embeddings from geographically proximate regions along similar latitudes exhibit stronger alignment. This spatial coherence indicates that geographically proximate areas exhibit systematically stronger cross-modal alignment.
@@ 5 Discussion & Conclusion
Limitations of our study include reliance on DHS clusters, which may introduce sampling biases, as remote areas are excluded from sampling (Kakooei and Daoud, 2024). The search agent may also retrieve post-treatment information from the web, potentially biasing causal analysis. Also, the computational demands of agent-driven pipelines at continental scales limit scalability for global applications.
Future work could integrate causal inference into multimodal AI, addressing text-specific challenges (Grimmer et al., 2022; Daoud et al., 2022; Pieuchon et al., 2024)—such as improved detection and control of post-treatment bias in agent representations—while further testing the Platonic Representation and Agent-Induced Novelty Hypotheses in open-ended contexts. $\otimes$